# PALM: A Dataset and Baseline for Learning Multi-subject Hand Prior

### v2

@@ Abstract
The ability to grasp objects, signal with gestures, and share emotion through touch all stem from the unique capabilities of human hands. Yet creating high-quality personalized hand avatars from images remains challenging due to complex geometry, appearance, and articulation, particularly under unconstrained lighting and limited views. Progress has also been limited by the lack of datasets that jointly provide accurate 3D geometry, high-resolution multi-view imagery, and a diverse population of subjects. To address this, we present PALM, a large-scale dataset comprising $13k$ high-quality hand scans from 263 subjects and $90k$ multi-view images, capturing rich variation in skin tone, age, and geometry. To show its utility, we present a baseline PALM-Net, a multi-subject prior over hand geometry and material properties learned via physically based inverse rendering, enabling realistic, relightable single-image hand avatar personalization. PALM’s scale and diversity make it a valuable real-world resource for hand modeling and related research.
@@ 1 Introduction
Human hands are central to how we interact with the physical and social world: we manipulate objects [hasson2019obman, wang2025magichoi, zhang2024artigrasp, fan2024benchmarks, fan2024hold], express intent through gestures [yu2025dynhamr, fan2021learning, tse2023spectral], and communicate affective cues via touch. Realistic and drivable hand avatars have the potential to transform virtual interaction, gaming, and telepresence. However, building such avatars from images remains a fundamentally challenging problem due to the complexity of hand geometry, appearance, and articulation, particularly under unconstrained lighting and from limited visual observations.
A critical missing component in the hand community is a large-scale, high-quality dataset that enables learning generalizable and physically grounded models of human hands. Existing datasets suffer from significant limitations: they often include only a small number of subjects [interhand, martinez2024codec], lack accurate 3D hand geometry from real-world scans [interhand, mano], or are derived from hand-crafted synthetic data [gao2022dart], limiting their utility for learning models that generalize across identity and illumination.
To fill this gap, we introduce PALM, a large-scale dataset of human hands containing publicly available-ready accurate hand scans, diverse in quantity and subject diversity. PALM includes $13k$ high-quality 3D hand scans and $90k$ high-resolution multi-view RGB images from $263$ subjects, each performing a diverse set of predefined hand poses designed to span a wide range of natural hand articulations. The subjects cover a broad range of skin tones and age groups. All data is captured using a commercial 3dMD scanner [3dmdhand], providing precise, sub-millimeter geometry. Each scan is paired with synchronized multi-view images and a MANO registration (pose and shape), obtained via multi-view–consistent alignment to the 3D scans. Importantly, the capture environment, including lighting and scanner configuration, remained fixed throughout the entire collection process, enabling consistent illumination conditions across subjects. While prior datasets have included either limited subjects or unreleased scan data, PALM will be made publicly available for research use upon publication, making it the most comprehensive and accessible dataset for studying generalizable hand models.
@@ 2 Related Work
Hand datasets: Recent years have seen an explosion of hand datasets, that can be categorized into hand-only [zimmermann2019freihand, moon2020deephandmesh, potamias2023handy, gao2022dart, martinez2024codec], interacting hand [tzionas2016capturing, interhand, moon2023dataset], and hand-object [hampali2020honnotate, fan2023arctic, dexycb, liu2022hoi4d, hasson2019obman, FirstPersonAction_CVPR2018, kwon2021h2o, contactpose_2020, sener2022assembly101, zhang2024graspxl] datasets. Early research on hand-only capture primarily focused on datasets collected using depth cameras [yuan2017bighand2]. Soon after, RGB-based datasets gained traction, with notable examples including STB [zhang20163d] and FreiHAND [zimmermann2019freihand]. Following this, interest expanded toward hand-object interaction datasets. Hampali et al[hampali2020honnotate] released a dataset of single hands manipulating rigid YCB objects, while Fan et al[fan2023arctic] used a marker-based MoCap setup to capture full-body interactions with articulated objects. More recently, Banerjee et al[banerjee2025hot3d] contributed a large-scale dataset featuring multi-object interactions. Interacting-hand datasets have also grown in popularity. Tzionas et al[tzionas2016capturing] pioneered capturing two-hand interactions using an RGB-D setup. Building on this, Moon et al[interhand] leveraged a large-scale multi-view RGB system to recover 3D hand poses under strong self-contact, which significantly boosted interest in modeling two interacting hands from RGB data. Moon et al[moon2023dataset] released a synthetic dataset using 3D annotations derived from [interhand]. Handy [potamias2023handy] provides 3dMD data but it is not publicly available. Despite the rapid emergence of various hand datasets, most approaches for learning hand avatars still rely primarily on InterHand2.6M [mundra2023livehand, chen2023handavatar] or on custom video recordings [karunratanakul2023harp]. This is largely due to the absence of a high-quality dataset that simultaneously provides accurate 3D hand scans, high-resolution multi-view RGB imagery, and a diverse set of subjects.
@@
Hand representations: Learning articulated hand representations is a long-standing research problem. Some methods solely focus on modelling hand joints [fan2021learning, spurr2020eccv, zimmermann2017iccv] or geometries [huangPhrit, moon2020deephandmesh, mano, duran2024hmp, ziani2022tempclr]. For example, MANO [mano] pioneered parametric mesh-based hand geometry modeling, parameterizing shape with a PCA latent space. Moon et al[moon2024authentic, moon2020deephandmesh] propose a non-linear approach for high-fidelity hand mesh modeling. HALO [Korrawe2021halo] is an implicit articulated hand geometry representation using occupancy network via a differentiable canonicalization layer. There are also methods that model both geometry and appearances of hands [potamias2023handy, qian2020html, chen2024urhand, li2022nimble, zheng2024ohta, corona2022lisa, mundra2023livehand, chen2023handavatar, fan2024hold]. HTML [qian2020html] extends MANO with a PCA-based texture model. Handy [potamias2023handy] is a parametric hand model of shape and texture learned from proprietary hand scans. Both Handy and HTML preprocess texture maps to minimize baked-in shadows and specularities. LISA and OHTA [corona2022lisa] model shape and appearance fields, with lighting effects baked into the network and controlled by latent codes. NIMBLE [li2022nimble] models hands with bone, muscle, and skin deformation, using a light stage to capture pose-independent albedo and specular maps modeled with PCA bases. HARP [karunratanakul2023harp] optimizes the normal and albedo maps for the MANO hand mesh with a point light source to model shadow effects, demonstrating slight generalizability to novel illuminations. URHand[chen2024urhand] models pose-dependent hand material properties and is trained on large-scale light stage data. Capturing such pose-dependent material characteristics and lighting effects requires accurate environment map information, typically enabled by a sophisticated light stage setup with hundreds of synchronized cameras, as used in Nimble and URHand. In contrast to these methods [corona2022lisa, zheng2024ohta, chen2024urhand], our baseline method can jointly learn appearances, geometries and relight the hand avatar and does not rely on expensive light stage setup.
@@ 3 PALMDataset
Overview: To study hand priors, we introduce PALM (see Figure 1), a high-quality dataset with accurate 3D hand annotations, high-resolution multi-view RGB images, and 3dMD hand scans. It contains $90k$ RGB images and $13k$ hand scans from 263 subjects (131 male, 132 female) with diverse skin tones and hand shapes. Data was captured using a 3dMD hand scanner [3dmdhand] with 7 RGB and 14 monochrome machine vision cameras calibrated; the RGB images have a resolution of $2448\times 2048$; and the 3D hand scans were reconstructed using 3dMD’s software-driven triangulation technique based on Active Stereo Photogrammetry. Participants performed approximately 50 predefined right-hand gestures. See SupMat for more examples in PALM.
@@ 3.1 Data Characteristics
ICVL [tang2014latent] 10 0 $320\times 240$ - track $\times$
BigHand2.2M [yuan2017bighand2] 10 0 $640\times 480$ - marker $\times$
Tzionas et al. [tzionas2016capturing] - 0 $640\times 480$ $\checkmark$ track $\times$
Simon et al. [simon2017hand] - 0 $1920\times 1080$ $\checkmark$ semi-auto $\times$
EgoDexter [mueller2017real] 4 0 $640\times 480$ $\checkmark$ manual $\times$
STB [zhang20163d] 1 0 $640\times 480$ $\checkmark$ manual $\times$
FreiHAND [Freihand2019] 32 0 $224\times 224$ $\checkmark$ semi-auto $\times$
InterHand2.6M [interhand] 50 0 $512\times 334$ $\checkmark$ semi-auto $\checkmark$
Re:InterHand [moon2023dataset] 10 0 $4096\times 2668$ $\times$ - $\times$
DART [gao2022dart] - 0 $512\times 512$ $\times$ - $\times$
MANO [mano] 31 1K N/A $\checkmark$ manual $\checkmark$
@@
Dataset comparison: Table 1 compares publicly available hand-only datasets. Early datasets [yuan2017bighand2, tzionas2016capturing, mueller2017real] focus on depth cameras and often have limited subjects and scale [tzionas2016capturing, mueller2017real]. Recent larger datasets focus on RGB images: InterHand2.6M [interhand] has 50 subjects with multi-view calibrated RGB, Re:InterHand [moon2023dataset] offers synthetically relit interacting-hand images of 10 subjects, and MANO [mano] provides hand scans of 31 subjects. Handy [potamias2023handy] includes 3dMD scans, but these are unavailable. While MANO, InterHand2.6M, and our dataset are all suitable for learning priors for hands, most other datasets are suboptimal due to missing modalities, limited subject diversity, synthetic data, or low image quality, leading most methods [moon2024authentic, mundra2023livehand, chen2023handavatar, karunratanakul2023harp] to rely on InterHand2.6M despite lacking scans. Our dataset is substantially larger in scale, with $263$ subjects, real high-resolution calibrated RGB images, and $13k$ high-quality scans, making it ideally suited for learning robust hand priors; it will be released publicly to advance future research.
@@ 3.2 Data Acquisition
Capture setup: We use a 3dMD [3dmdhand] hand scanner to capture high-resolution multi-view RGB images and 3D scans of the subjects. In particular, the 7-viewpoint setup consists of an array of 21 synchronized and calibrated machine vision cameras, including both RGB and monochrome sensors, arranged to provide full 360-degree coverage of the hand. The system captures images at a high resolution of $2448\times 2048$. Random light projectors are integrated to enhance surface detail and geometry accuracy. The 3D hand scans were reconstructed using 3dMD’s software-driven triangulation technique based on Active Stereo Photogrammetry. All cameras are calibrated, ensuring consistent alignment across views. This setup enables the acquisition of dense, accurate 3D hand meshes in various poses, suitable for studying both static poses and dynamic hand articulations.
@@
3D keypoint annotation: A semi-automatic pipeline is used to generate accurate 3D hand pose labels. The 2D keypoints are first manually annotated on a subset of images to train a 2D keypoint detector tailored to our capture setup. The detector, implemented as a U-Net [ronneberger2015u] pre-trained on InterHand2.6M and fine-tuned on 10K manually annotated PALM images, is then used to estimate 2D keypoints for all camera views, which are subsequently triangulated to obtain 3D poses using multi-view geometry. Specifically, we follow the approach used in InterHand2.6M and apply RANSAC-based triangulation to robustly solve for the 3D keypoint locations. This semi-automatic approach significantly reduces the manual labeling burden while ensuring reliable 3D pose annotations.
MANO registration: The 3D MANO poses for each pose and subject are obtained by registering the MANO hand model to the hand scans. To register MANO to each hand gesture, we optimize the MANO hand model using a combination of 2D/3D keypoints, segmentation mask, and 3D hand scan supervision. Specifically, we minimize the closest-point distance from each MANO vertex to the corresponding hand scan surface. Ground-truth masks are derived from the hand scans, and we employ Soft Rasterizer [liu2019soft] for differentiable rendering to align the MANO silhouette with these masks.
The registration is independently performed for each subject in two stages. The first stage optimizes the per-subject hand shape parameters and per-frame pose parameters on a set of simple hand poses (e.g., flat hand), and the second stage only optimizes the per-frame pose parameters while freezing the shape parameters. Our final 3D keypoints have a recall rate of 95% at 10mm threshold. Our MANO registrations show a mean fitting error of 5.3 mm with respect to the 3D keypoints (similar to that of InterHand2.6M [interhand]).
@@ 4 Method
PALM-Net, illustrated in Figure 4, utilizes PALM, our large-scale collection of human hand data containing detailed hand scans and high-resolution RGB images, to train a personalizable and relightable hand prior. In this section, we first introduce preliminary concepts on NeRF [mildenhall2021nerf], the Neural Radiance Field technique at the core of our approach, and MANO  [mano], the parametric hand model that we leverage within our pipeline to map 3D points across different subjects and hand poses into a shared canonical representation. Next, we present PALM-Net, our novel framework for learning a multi-subject hand prior through physically based inverse rendering (PBR), and describe how to train our representation over multiple subjects. Finally, we detail how PALM-Net can be used to recover a personalized and relightable hand avatar from a single image. This is a highly under-constrained problem, and we show that a prior model such as PALM-Net helps in recovering realistic, personalized hand avatars even in extreme illumination settings.
@@ 4.1 Preliminaries
NeRF: Given a ray $\mathbf{r}=(\mathbf{o},\mathbf{d})$ defined by its camera center $\mathbf{o}$ and viewing direction $\mathbf{d}$, NeRF [mildenhall2021nerf] computes the output radiance (i.e., pixel color) of the ray via:
@@
The MANO  [mano] hand model is parametrized by $\mathbf{\Theta}=\{\mathbf{\theta},\mathbf{\beta},\mathbf{p}\}$, where $\mathbf{\theta}\in\rm I\!R^{45}$ denotes hand skeletal pose (joint angles), $\mathbf{\beta}\in\rm I\!R^{10}$ hand shape (parameterized by PCA coefficients) and $\mathbf{p}\in\rm I\!R^{6}$ global transformation. The MANO model then maps $\mathbf{\Theta}$ to a posed 3D mesh $\mathcal{M}(\mathbf{\Theta})\in\rm I\!R^{778\times 3}$. PALM-Net leverages MANO to perform inverse LBS using SNARF [chen2021snarf] that map points in 3D to a common canonical representation, i.e., given a 3D point in deformed coordinates $\mathbf{x}_{d}$, hand parameters $\mathbf{\Theta}$, we find the corresponding 3D point $\mathbf{x}_{c}$ in canonical space as,

### v3

@@ Abstract
The ability to grasp objects, signal with gestures, and share emotion through touch all stem from the unique capabilities of human hands. Yet creating high-quality personalized hand avatars from images remains challenging due to complex geometry, appearance, and articulation, particularly under unconstrained lighting and limited views. Progress has also been limited by the lack of datasets that jointly provide accurate 3D geometry, high-resolution multi-view imagery, and a diverse population of subjects. To address this, we present PALM, a large-scale dataset comprising $13k$ high-quality hand scans from 263 subjects and $90k$ multi-view images, capturing rich variation in skin tone, age, and geometry. To show its utility, we present a baseline PALM-Net, a multi-subject prior over hand geometry and material properties learned via physically based inverse rendering, enabling realistic, relightable single-image hand avatar personalization. PALM’s scale and diversity make it a valuable real-world resource for hand modeling and related research. See the project page at this link.
$\dagger$ Work done during Meta internship
@@ 1 Introduction
Human hands are central to how we interact with the physical and social world: we manipulate objects [19, 47, 52, 11, 12], express intent through gestures [49, 13, 44], and communicate affective cues via touch. Realistic and drivable hand avatars have the potential to transform virtual interaction, gaming, and telepresence. However, building such avatars from images remains a fundamentally challenging problem due to the complexity of hand geometry, appearance, and articulation, particularly under unconstrained lighting and from limited visual observations.
A critical missing component in the hand community is a large-scale, high-quality dataset that enables learning generalizable and physically grounded models of human hands. Existing datasets suffer from significant limitations: they often include only a small number of subjects [32, 27], lack accurate 3D hand geometry from real-world scans [32, 37], or are derived from hand-crafted synthetic data [15], limiting their utility for learning models that generalize across identity and illumination.
To fill this gap, we introduce PALM, a large-scale dataset of human hands containing publicly available-ready accurate hand scans, diverse in quantity and subject diversity. PALM includes $13k$ high-quality 3D hand scans and $90k$ high-resolution multi-view RGB images from $263$ subjects, each performing a diverse set of predefined hand poses designed to span a wide range of natural hand articulations. The subjects cover a broad range of skin tones and age groups. All data is captured using a commercial 3dMD scanner [43], providing precise, sub-millimeter geometry. Each scan is paired with synchronized multi-view images and a MANO registration (pose and shape), obtained via multi-view–consistent alignment to the 3D scans. Importantly, the capture environment, including lighting and scanner configuration, remained fixed throughout the entire collection process, enabling consistent illumination conditions across subjects. While prior datasets have included either limited subjects or unreleased scan data, PALM will be made publicly available for research use upon publication, making it the most comprehensive and accessible dataset for studying generalizable hand models.
@@ 2 Related Work
Hand datasets: Recent years have seen an explosion of hand datasets, that can be categorized into hand-only [59, 30, 35, 15, 27], interacting hand [45, 32, 29], and hand-object [18, 14, 4, 26, 19, 16, 23, 2, 39, 51] datasets. Early research on hand-only capture primarily focused on datasets collected using depth cameras [50]. Soon after, RGB-based datasets gained traction, with notable examples including STB [53] and FreiHAND [59]. Following this, interest expanded toward hand-object interaction datasets. Hampali et al. [18] released a dataset of single hands manipulating rigid YCB objects, while Fan et al. [14] used a marker-based MoCap setup to capture full-body interactions with articulated objects. More recently, Banerjee et al. [1] contributed a large-scale dataset featuring multi-object interactions. Interacting-hand datasets have also grown in popularity. Tzionas et al. [45] pioneered capturing two-hand interactions using an RGB-D setup. Building on this, Moon et al. [32] leveraged a large-scale multi-view RGB system to recover 3D hand poses under strong self-contact, which significantly boosted interest in modeling two interacting hands from RGB data. Moon et al. [29] released a synthetic dataset using 3D annotations derived from [32]. Handy [35] provides 3dMD data but it is not publicly available. Despite the rapid emergence of various hand datasets, most approaches for learning hand avatars still rely primarily on InterHand2.6M [34, 5] or on custom video recordings [21]. This is largely due to the absence of a high-quality dataset that simultaneously provides accurate 3D hand scans, high-resolution multi-view RGB imagery, and a diverse set of subjects.
@@
Hand representations: Learning articulated hand representations is a long-standing research problem. Some methods solely focus on modelling hand joints [13, 41, 57] or geometries [20, 30, 37, 10, 56]. For example, MANO [37] pioneered parametric mesh-based hand geometry modeling, parameterizing shape with a PCA latent space. Moon et al. [31, 30] propose a non-linear approach for high-fidelity hand mesh modeling. HALO [22] is an implicit articulated hand geometry representation using occupancy network via a differentiable canonicalization layer. There are also methods that model both geometry and appearances of hands [35, 36, 8, 24, 55, 9, 34, 5, 12]. HTML [36] extends MANO with a PCA-based texture model. Handy [35] is a parametric hand model of shape and texture learned from proprietary hand scans. Both Handy and HTML preprocess texture maps to minimize baked-in shadows and specularities. LISA and OHTA [9] model shape and appearance fields, with lighting effects baked into the network and controlled by latent codes. NIMBLE [24] models hands with bone, muscle, and skin deformation, using a light stage to capture pose-independent albedo and specular maps modeled with PCA bases. HARP [21] optimizes the normal and albedo maps for the MANO hand mesh with a point light source to model shadow effects, demonstrating slight generalizability to novel illuminations. URHand[8] models pose-dependent hand material properties and is trained on large-scale light stage data. Capturing such pose-dependent material characteristics and lighting effects requires accurate environment map information, typically enabled by a sophisticated light stage setup with hundreds of synchronized cameras, as used in Nimble and URHand. In contrast to these methods [9, 55, 8], our baseline method can jointly learn appearances, geometries and relight the hand avatar and does not rely on expensive light stage setup.
@@ 3 PALMDataset
Overview: To study hand priors, we introduce PALM (see Figure 1), a high-quality dataset with accurate 3D hand annotations, high-resolution multi-view RGB images, and 3dMD hand scans. It contains $90k$ RGB images and $13k$ hand scans from 263 subjects (131 male, 132 female) with diverse skin tones and hand shapes. Data was captured using a 3dMD hand scanner [43] with 7 RGB and 14 monochrome machine vision cameras calibrated; the RGB images have a resolution of $2448\times 2048$; and the 3D hand scans were reconstructed using 3dMD’s software-driven triangulation technique based on Active Stereo Photogrammetry. Participants performed approximately 50 predefined right-hand gestures. See SupMat for more examples in PALM.
@@ 3.1 Data Characteristics
ICVL [42] 10 0 $320\times 240$ - track $\times$
BigHand2.2M [50] 10 0 $640\times 480$ - marker $\times$
Tzionas et al. [45] - 0 $640\times 480$ $\checkmark$ track $\times$
Simon et al. [40] - 0 $1920\times 1080$ $\checkmark$ semi-auto $\times$
EgoDexter [33] 4 0 $640\times 480$ $\checkmark$ manual $\times$
STB [53] 1 0 $640\times 480$ $\checkmark$ manual $\times$
FreiHAND [58] 32 0 $224\times 224$ $\checkmark$ semi-auto $\times$
InterHand2.6M [32] 50 0 $512\times 334$ $\checkmark$ semi-auto $\checkmark$
Re:InterHand [29] 10 0 $4096\times 2668$ $\times$ - $\times$
DART [15] - 0 $512\times 512$ $\times$ - $\times$
MANO [37] 31 1K N/A $\checkmark$ manual $\checkmark$
@@
Dataset comparison: Table 1 compares publicly available hand-only datasets. Early datasets [50, 45, 33] focus on depth cameras and often have limited subjects and scale [45, 33]. Recent larger datasets focus on RGB images: InterHand2.6M [32] has 50 subjects with multi-view calibrated RGB, Re:InterHand [29] offers synthetically relit interacting-hand images of 10 subjects, and MANO [37] provides hand scans of 31 subjects. Handy [35] includes 3dMD scans, but these are unavailable. While MANO, InterHand2.6M, and our dataset are all suitable for learning priors for hands, most other datasets are suboptimal due to missing modalities, limited subject diversity, synthetic data, or low image quality, leading most methods [31, 34, 5, 21] to rely on InterHand2.6M despite lacking scans. Our dataset is substantially larger in scale, with $263$ subjects, real high-resolution calibrated RGB images, and $13k$ high-quality scans, making it ideally suited for learning robust hand priors; it will be released publicly to advance future research.
@@ 3.2 Data Acquisition
Capture setup: We use a 3dMD [43] hand scanner to capture high-resolution multi-view RGB images and 3D scans of the subjects. In particular, the 7-viewpoint setup consists of an array of 21 synchronized and calibrated machine vision cameras, including both RGB and monochrome sensors, arranged to provide full 360-degree coverage of the hand. The system captures images at a high resolution of $2448\times 2048$. Random light projectors are integrated to enhance surface detail and geometry accuracy. The 3D hand scans were reconstructed using 3dMD’s software-driven triangulation technique based on Active Stereo Photogrammetry. All cameras are calibrated, ensuring consistent alignment across views. This setup enables the acquisition of dense, accurate 3D hand meshes in various poses, suitable for studying both static poses and dynamic hand articulations.
@@
3D keypoint annotation: A semi-automatic pipeline is used to generate accurate 3D hand pose labels. The 2D keypoints are first manually annotated on a subset of images to train a 2D keypoint detector tailored to our capture setup. The detector, implemented as a U-Net [38] pre-trained on InterHand2.6M and fine-tuned on 10K manually annotated PALM images, is then used to estimate 2D keypoints for all camera views, which are subsequently triangulated to obtain 3D poses using multi-view geometry. Specifically, we follow the approach used in InterHand2.6M and apply RANSAC-based triangulation to robustly solve for the 3D keypoint locations. This semi-automatic approach significantly reduces the manual labeling burden while ensuring reliable 3D pose annotations.
MANO registration: The 3D MANO poses for each pose and subject are obtained by registering the MANO hand model to the hand scans. To register MANO to each hand gesture, we optimize the MANO hand model using a combination of 2D/3D keypoints, segmentation mask, and 3D hand scan supervision. Specifically, we minimize the closest-point distance from each MANO vertex to the corresponding hand scan surface. Ground-truth masks are derived from the hand scans, and we employ Soft Rasterizer [25] for differentiable rendering to align the MANO silhouette with these masks.
The registration is independently performed for each subject in two stages. The first stage optimizes the per-subject hand shape parameters and per-frame pose parameters on a set of simple hand poses (e.g., flat hand), and the second stage only optimizes the per-frame pose parameters while freezing the shape parameters. Our final 3D keypoints have a recall rate of 95% at 10mm threshold. Our MANO registrations show a mean fitting error of 5.3 mm with respect to the 3D keypoints (similar to that of InterHand2.6M [32]).
@@ 4 Method
PALM-Net, illustrated in Figure 4, utilizes PALM, our large-scale collection of human hand data containing detailed hand scans and high-resolution RGB images, to train a personalizable and relightable hand prior. In this section, we first introduce preliminary concepts on NeRF [28], the Neural Radiance Field technique at the core of our approach, and MANO  [37], the parametric hand model that we leverage within our pipeline to map 3D points across different subjects and hand poses into a shared canonical representation. Next, we present PALM-Net, our novel framework for learning a multi-subject hand prior through physically based inverse rendering (PBR), and describe how to train our representation over multiple subjects. Finally, we detail how PALM-Net can be used to recover a personalized and relightable hand avatar from a single image. This is a highly under-constrained problem, and we show that a prior model such as PALM-Net helps in recovering realistic, personalized hand avatars even in extreme illumination settings.
@@ 4.1 Preliminaries
NeRF: Given a ray $\mathbf{r}=(\mathbf{o},\mathbf{d})$ defined by its camera center $\mathbf{o}$ and viewing direction $\mathbf{d}$, NeRF [28] computes the output radiance (i.e., pixel color) of the ray via:
@@
The MANO  [37] hand model is parametrized by $\mathbf{\Theta}=\{\mathbf{\theta},\mathbf{\beta},\mathbf{p}\}$, where $\mathbf{\theta}\in\rm I\!R^{45}$ denotes hand skeletal pose (joint angles), $\mathbf{\beta}\in\rm I\!R^{10}$ hand shape (parameterized by PCA coefficients) and $\mathbf{p}\in\rm I\!R^{6}$ global transformation. The MANO model then maps $\mathbf{\Theta}$ to a posed 3D mesh $\mathcal{M}(\mathbf{\Theta})\in\rm I\!R^{778\times 3}$. PALM-Net leverages MANO to perform inverse LBS using SNARF [7] that map points in 3D to a common canonical representation, i.e., given a 3D point in deformed coordinates $\mathbf{x}_{d}$, hand parameters $\mathbf{\Theta}$, we find the corresponding 3D point $\mathbf{x}_{c}$ in canonical space as,